OpenStudy (anonymous):

Please help me solve! :) 12C5= 12!/7!5!

OpenStudy (anonymous):

@whpalmer4 hi! do you think you could help me out? :)

OpenStudy (solomonzelman):

\(\Huge\color{royalblue}{ \sf 12~~C~~5= \frac{12!}{5!~\times ~(12-5)i}=\frac{12!}{5!~\times ~7i} }\) yes you got the formula right, so I am not sure what exactly did you need? perhaps you just wanted the calculator? \(\large\color{blue}{ \cal combinations }\) http://www.calculatorsoup.com/calculators/discretemathematics/combinations.php \(\large\color{blue}{ \cal permutations }\) http://www.calculatorsoup.com/calculators/discretemathematics/permutations.php

OpenStudy (anonymous):

Fill in the details \[ \frac{12!}{5! 7!}=\frac{8\times 9\times 10\times 11\times12}{2\times 3\times 4\times 5}=8\times 3\times 11\times 3=792 \]